# SHAMISA: SHAped Modeling of Implicit Structural Associations for Self-supervised No-Reference Image Quality Assessment

### 14 Mar 2026 v1

@@ I Introduction
At the core of SHAMISA is a compositional distortion engine that generates infinite compositional degradations from continuous parameter spaces. Each mini-batch is built from reference images that form distortion composition groups where only one degradation factor varies, enabling controlled sampling across content, distortion type, and severity. From this, we build two categories of relation graphs: (i) Metadata-Driven Graphs, which encode pairwise similarity based on distortion metadata, encouraging images with similar degradations to lie close in the learned manifold while inducing controlled representational shifts as distortion severity varies; and (ii) Structurally Intrinsic Graphs, constructed from the latent feature space using $k$-nearest neighbors (kNN) and deep clustering. These graphs supervise a non-contrastive VICReg-style objective, optimized in a single update that constructs relation graphs from the current representations and applies stop-gradient through graph construction.
@@ IV-A Setup
We pre-train for one epoch over the 140k KADIS reference images with SGD. Encoder and projector dimensions are $d_{h}{=}2048$ and $d_{z}{=}256$. Each update uses B=2B{=}2 tiny-batches, R=3R{=}3 references, C=4C{=}4 composition groups, and L=5L{=}5 severity levels, giving $N_{\text{ref}}{=}BR{=}6$, $N_{\text{comp}}{=}CL{=}20$, and $N{=}N_{\text{ref}}(N_{\text{comp}}+1){=}126$ in Eq. 1. Distortion generation uses $M_{d}{=}7$. Graph sparsity and OT settings are $k_{n}{=}31$, $K_{d}{=}4096$, K=32K{=}32, and $\tau_{c}{=}0.1$; for the OT graph, the effective global sparsification budget scales with mini-batch size through $K_{g}{=}8N$. Crop size is $224\times 224$. Remaining details are listed in §A.
@@ IV-B Datasets and Evaluation Protocol
We include both synthetic and in-the-wild benchmarks. Synthetic: LIVE [41], CSIQ [42], TID2013 [43], KADID-10K [9]. In-the-wild: FLIVE [29], SPAQ [44]. Brief dataset statistics are available in Appendix D.
@@ IV-E Cross-Dataset Transfer
We follow the standard SSL-IQA linear-probe setting: the encoder is frozen; we train a ridge regressor on the train split of the source dataset, select its regularization on the source val split, and evaluate zero-shot on the target test split with identical multi-scale features and five-crop pooling. We report the median over 10 reference-disjoint splits. This transfer style mirrors practical deployment. A large pool of unlabeled images is used for SSL pre-training, and only a small labeled set from a related source dataset is available to fit a lightweight head. The model must then generalize to a related but distinct target without target labels and without adapting normalization statistics to the target dataset. Full details are provided in Appendix F.
@@
Figure 2: gMAD on Waterloo Exploration (distorted pool): panels (a,b) use SHAMISA as the defender and panels (c,d) use ARNIQA as the defender. In each panel, the top and bottom images are the attacker-selected pair from the indicated defender-quality bin, with low-bin cases shown in (a,c) and high-bin cases shown in (b,d). Appendix Table VII reports the corresponding top-1 attacker gaps.
We use the group maximum differentiation (gMAD) protocol [48] as a pairwise diagnostic on the Waterloo Exploration Database [49] using its distorted image pool (no MOS annotations). Following our standard NR-IQA evaluation pipeline, we convert frozen representations into scalar quality predictions by fitting a ridge regressor on KADID-10K and then apply the resulting predictors to all Waterloo images. Given a defender model, we first split the pool into equal-count bins based on the defender’s predicted quality. Within a chosen bin, the attacker then selects the image pair with the largest attacker-side score gap while the defender still scores the pair similarly. Appendix I gives the corresponding protocol details. Fig. 2 shows representative pairs for low and high defender levels in both defender-attacker directions, and Appendix Table VII reports the corresponding top-1 attacker gaps. In this pairwise Waterloo gMAD comparison with KADID-trained ridge heads, attacker gaps are smaller when SHAMISA is the defender than when ARNIQA is the defender, suggesting tighter within-bin quality consistency for SHAMISA in this specific setting.
@@
We visualize the structure of the learned representation space by applying 2D t-SNE [50] to image-level encoder representations $\mathbf{H}$ on KADID-10K [9]. For each image, we extract $\mathbf{H}$ (pre-projector) and average across crops, yielding 10,12510{,}125 points. Fig. 3 colors points by coarse distortion family using a shared, fixed palette for fair visual comparison. Both methods separate several major degradation families, but SHAMISA often forms more coherent regions for prominent categories, whereas ARNIQA shows more inter-family mixing in the central area. We provide additional severity-colored visualizations and subtype-level views (by distortion family) in Appendix G. Across these plots, separability is strongest for families with distinctive artifact statistics (e.g., blur, noise, and compression), while brightness-related manipulations remain more entangled and challenging for both models. Overall, these plots are consistent with SHAMISA’s pre-training encouraging a representation geometry that is coherent across major distortion families while still preserving fine-grained variation within them.
@@ IV-I Ablations
TABLE III: Main ablations. Median over 10 splits per dataset; then averaged. SRCCs/PLCCs average over synthetic $\{\mathrm{LIVE},\mathrm{CSIQ},\mathrm{TID2013},\mathrm{KADID\mbox{-}10K}\}$, SRCCa/PLCCa over authentic $\{\mathrm{FLIVE},\mathrm{SPAQ}\}$, and SRCCg/PLCCg denote the average over the six per-dataset median SRCC/PLCC values. Full ablation catalog is reported in Appendix Table VIII.
@@
We ablate SHAMISA along complementary axes that map directly to our main design choices: (A) how we aggregate multi-source relations into the soft matrix $\bm{G}$} (learned end-to-end mixing $\Phi$ vs fixed mixing; regularization), (B) which relation sources are necessary (structure graphs ${\bm{G}}^{k},{\bm{G}}^{o}$ vs metadata graphs ${\bm{G}}^{rd},{\bm{G}}^{dd},{\bm{G}}^{rr}$), (C) how we sparsify the global OT graph (global vs row-wise truncation), (D) how the compositional distortion engine defines the self-supervised training distribution (composition complexity, severity modeling, ordering), and (E–F) objective diagnostics that isolate the contribution of OT alignment and severity-dependent edge weighting. Each variant changes only minimal factors relative to the reference configuration (A0). We evaluate on six NR-IQA datasets (LIVE, CSIQ, TID2013, KADID-10K, FLIVE, SPAQ) with 10 random splits per dataset, reporting the median test SRCC/PLCC per dataset, then summarizing synthetic and authentic groups separately together with a global comparison score. For readability, Table III highlights the most diagnostic variants; the complete catalog is deferred to Appendix Table VIII.
@@ IV-J Pre-training Dynamics
Importantly, the best checkpoint can be dataset-dependent: while some datasets continue to improve or remain stable as pre-training proceeds, others exhibit a mild peak-and-decline behavior, which is expected because the SSL objective is not optimized directly for SRCC on any single benchmark. As a deployment note, when a small labeled validation set is available in the target domain, it can be used to select a checkpoint for that domain. Unless stated otherwise, all results reported in Tables I and II use the fixed A0 checkpoint; this section is a diagnostic and is not used for model selection. Additional diagnostic figures are provided in Appendix K.
@@ IV-K Hyperparameter Sensitivity
We analyze hyperparameter sensitivity by varying one knob at a time around our A0 configuration while keeping the rest fixed. Fig. 6 shows that most capacity and graph-construction choices admit broad plateaus: $d_{h}$, $d_{z}$, and $k_{n}$ typically yield stable SRCC across a wide range, and increasing $K$ mainly saturates after moderate values. In contrast, the OT alignment coefficient $\eta$ is the most influential knob across datasets, with the largest SRCC swings on synthetic distortion benchmarks such as TID2013 and KADID-10K, suggesting that overly weak alignment under-utilizes relational guidance, whereas overly strong alignment can over-constrain representations. Among VICReg terms, the invariance weight $\gamma$ exhibits the clearest sensitivity, particularly on FLIVE, whereas the variance and covariance weights $\alpha$ and $\beta$ are comparatively robust across the tested ranges. Overall, these sweeps indicate that SHAMISA is robust across broad ranges of capacity and graph-design choices, while a smaller set of alignment-related coefficients, especially $\eta$ and $\gamma$, remains important for fine control. Appendix L reports the corresponding SRCC sensitivity summary.
@@
We also propose a full-reference variant, SHAMISA-FR, that reuses the frozen no-reference encoder and uses the absolute feature difference $|h_{\mathrm{ref}}{-}h_{\mathrm{dist}}|$ between pristine and distorted images as the regressor input. This variant requires no additional backbone pre-training and attains competitive FR-IQA accuracy; details are provided in Appendix E, with the main comparison in Appendix Table VI.
@@
Additional implementation details, optimization settings, evaluation protocols, and supplementary diagnostics are provided in Appendices A, F, K, and L. The project repository is available at https://github.com/Mahdi-Naseri/SHAMISA/.

### 17 Mar 2026 v2

@@ I Introduction
At the core of SHAMISA is a compositional distortion engine that generates infinite compositional degradations from continuous parameter spaces. Each mini-batch is built from reference images that form distortion composition groups where only one degradation factor varies, enabling controlled sampling across content, distortion type, and severity. From this, we build two categories of relation graphs: (i) Metadata-Driven Graphs, which encode pairwise similarity based on distortion metadata, encouraging images with similar degradations to lie close in the learned manifold while inducing controlled representational shifts as distortion severity varies; and (ii) Structurally Intrinsic Graphs, constructed from the latent feature space using $k$-nearest neighbors (kNN) and deep clustering. These graphs supervise a non-contrastive VICReg-style objective with graph-weighted invariance. With stop-gradient applied to graph construction, each iteration builds relation graphs from the current representations and updates the model under the resulting objective in a single optimization step.
@@ IV-A Setup
We pre-train for one epoch over the 140k KADIS reference images with SGD. Encoder and projector dimensions are $d_{h}{=}2048$ and $d_{z}{=}256$. Each update uses B=2B{=}2 tiny-batches, R=3R{=}3 references, C=4C{=}4 composition groups, and L=5L{=}5 severity levels, giving $N_{\text{ref}}{=}BR{=}6$, $N_{\text{comp}}{=}CL{=}20$, and $N{=}N_{\text{ref}}(N_{\text{comp}}+1){=}126$ in Eq. 1. Distortion generation uses $M_{d}{=}7$. Graph sparsity and OT settings are $k_{n}{=}31$, $K_{d}{=}4096$, K=32K{=}32, and $\tau_{c}{=}0.1$; for the OT graph, the effective global sparsification budget scales with mini-batch size through $K_{g}{=}8N$. Crop size is $224\times 224$. Remaining details are listed in Supplementary Appendix A.
@@ IV-B Datasets and Evaluation Protocol
We include both synthetic and in-the-wild benchmarks. Synthetic: LIVE [41], CSIQ [42], TID2013 [43], KADID-10K [9]. In-the-wild: FLIVE [29], SPAQ [44]. Brief dataset statistics are available in Supplementary Appendix D.
@@ IV-E Cross-Dataset Transfer
We follow the standard SSL-IQA linear-probe setting: the encoder is frozen; we train a ridge regressor on the train split of the source dataset, select its regularization on the source val split, and evaluate zero-shot on the target test split with identical multi-scale features and five-crop pooling. We report the median over 10 reference-disjoint splits. This transfer style mirrors practical deployment. A large pool of unlabeled images is used for SSL pre-training, and only a small labeled set from a related source dataset is available to fit a lightweight head. The model must then generalize to a related but distinct target without target labels and without adapting normalization statistics to the target dataset. Full details are provided in Supplementary Appendix F.
@@
Figure 2: gMAD on Waterloo Exploration (distorted pool): panels (a,b) use SHAMISA as the defender and panels (c,d) use ARNIQA as the defender. In each panel, the top and bottom images are the attacker-selected pair from the indicated defender-quality bin, with low-bin cases shown in (a,c) and high-bin cases shown in (b,d). Supplementary Table VII reports the corresponding top-1 attacker gaps.
We use the group maximum differentiation (gMAD) protocol [48] as a pairwise diagnostic on the Waterloo Exploration Database [49] using its distorted image pool (no MOS annotations). Following our standard NR-IQA evaluation pipeline, we convert frozen representations into scalar quality predictions by fitting a ridge regressor on KADID-10K and then apply the resulting predictors to all Waterloo images. Given a defender model, we first split the pool into equal-count bins based on the defender’s predicted quality. Within a chosen bin, the attacker then selects the image pair with the largest attacker-side score gap while the defender still scores the pair similarly. Supplementary Appendix I gives the corresponding protocol details. Fig. 2 shows representative pairs for low and high defender levels in both defender-attacker directions, and Supplementary Table VII reports the corresponding top-1 attacker gaps. In this pairwise Waterloo gMAD comparison with KADID-trained ridge heads, attacker gaps are smaller when SHAMISA is the defender than when ARNIQA is the defender, suggesting tighter within-bin quality consistency for SHAMISA in this specific setting.
@@
We visualize the structure of the learned representation space by applying 2D t-SNE [50] to image-level encoder representations $\mathbf{H}$ on KADID-10K [9]. For each image, we extract $\mathbf{H}$ (pre-projector) and average across crops, yielding 10,12510{,}125 points. Fig. 3 colors points by coarse distortion family using a shared, fixed palette for fair visual comparison. Both methods separate several major degradation families, but SHAMISA often forms more coherent regions for prominent categories, whereas ARNIQA shows more inter-family mixing in the central area. We provide additional severity-colored visualizations and subtype-level views (by distortion family) in Supplementary Appendix G. Across these plots, separability is strongest for families with distinctive artifact statistics (e.g., blur, noise, and compression), while brightness-related manipulations remain more entangled and challenging for both models. Overall, these plots are consistent with SHAMISA’s pre-training encouraging a representation geometry that is coherent across major distortion families while still preserving fine-grained variation within them.
@@ IV-I Ablations
TABLE III: Main ablations. Median over 10 splits per dataset; then averaged. SRCCs/PLCCs average over synthetic $\{\mathrm{LIVE},\mathrm{CSIQ},\mathrm{TID2013},\mathrm{KADID\mbox{-}10K}\}$, SRCCa/PLCCa over authentic $\{\mathrm{FLIVE},\mathrm{SPAQ}\}$, and SRCCg/PLCCg denote the average over the six per-dataset median SRCC/PLCC values. Full ablation catalog is reported in Supplementary Table VIII.
@@
We ablate SHAMISA along complementary axes that map directly to our main design choices: (A) how we aggregate multi-source relations into the soft matrix $\bm{G}$} (learned end-to-end mixing $\Phi$ vs fixed mixing; regularization), (B) which relation sources are necessary (structure graphs ${\bm{G}}^{k},{\bm{G}}^{o}$ vs metadata graphs ${\bm{G}}^{rd},{\bm{G}}^{dd},{\bm{G}}^{rr}$), (C) how we sparsify the global OT graph (global vs row-wise truncation), (D) how the compositional distortion engine defines the self-supervised training distribution (composition complexity, severity modeling, ordering), and (E–F) objective diagnostics that isolate the contribution of OT alignment and severity-dependent edge weighting. Each variant changes only minimal factors relative to the reference configuration (A0). We evaluate on six NR-IQA datasets (LIVE, CSIQ, TID2013, KADID-10K, FLIVE, SPAQ) with 10 random splits per dataset, reporting the median test SRCC/PLCC per dataset, then summarizing synthetic and authentic groups separately together with a global comparison score. For readability, Table III highlights the most diagnostic variants; the complete catalog is deferred to Supplementary Table VIII.
@@ IV-J Pre-training Dynamics
Importantly, the best checkpoint can be dataset-dependent: while some datasets continue to improve or remain stable as pre-training proceeds, others exhibit a mild peak-and-decline behavior, which is expected because the SSL objective is not optimized directly for SRCC on any single benchmark. As a deployment note, when a small labeled validation set is available in the target domain, it can be used to select a checkpoint for that domain. Unless stated otherwise, all results reported in Tables I and II use the fixed A0 checkpoint; this section is a diagnostic and is not used for model selection. Additional diagnostic figures are provided in Supplementary Appendix K.
@@ IV-K Hyperparameter Sensitivity
We analyze hyperparameter sensitivity by varying one knob at a time around our A0 configuration while keeping the rest fixed. Fig. 6 shows that most capacity and graph-construction choices admit broad plateaus: $d_{h}$, $d_{z}$, and $k_{n}$ typically yield stable SRCC across a wide range, and increasing $K$ mainly saturates after moderate values. In contrast, the OT alignment coefficient $\eta$ is the most influential knob across datasets, with the largest SRCC swings on synthetic distortion benchmarks such as TID2013 and KADID-10K, suggesting that overly weak alignment under-utilizes relational guidance, whereas overly strong alignment can over-constrain representations. Among VICReg terms, the invariance weight $\gamma$ exhibits the clearest sensitivity, particularly on FLIVE, whereas the variance and covariance weights $\alpha$ and $\beta$ are comparatively robust across the tested ranges. Overall, these sweeps indicate that SHAMISA is robust across broad ranges of capacity and graph-design choices, while a smaller set of alignment-related coefficients, especially $\eta$ and $\gamma$, remains important for fine control. Supplementary Appendix L reports the corresponding SRCC sensitivity summary.
@@
We also propose a full-reference variant, SHAMISA-FR, that reuses the frozen no-reference encoder and uses the absolute feature difference $|h_{\mathrm{ref}}{-}h_{\mathrm{dist}}|$ between pristine and distorted images as the regressor input. This variant requires no additional backbone pre-training and attains competitive FR-IQA accuracy; details are provided in Supplementary Appendix E, with the main comparison in Supplementary Table VI.
@@
Additional implementation details, optimization settings, evaluation protocols, and supplementary diagnostics are provided in Supplementary Appendix A, Supplementary Appendix F, Supplementary Appendix K, and Supplementary Appendix L. The project repository is available at https://github.com/Mahdi-Naseri/SHAMISA/.